OpenStudy (anonymous):

I have to find the exact value of arccos(sin(pi/6)). How would I do that?

OpenStudy (anonymous):

That would be radians, not degrees, right? And how do I do it without a calculator?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

let arc cos(sin(pi/6))=x then cosx=sin(pi/6)=1/2=cos(pi/3), x=pi/3